1.12L of gas has mass 3g then find molecular mass of gas and V.D

Grade:11

1 Answers

Moh jaced
5 years ago
1.12L gas = 3 g
So 22.4L gas = (3/1.12)×22.4 = 60g
So the molecular wt is 60
But molecular = 2× VD
So V.D = 30